OpenStudy (anonymous):

The speed of a passenger train is 16mph faster than the speed of a freight train. The passenger train travels 300 miles in the same time it takes the freight train to travel 220 miles. Find the speed of each train.

OpenStudy (anonymous):

Call the speed of the freight train x. What's the speed of the passenger train? Using the formula distance = speed * time, how long does it take for the passenger train to go 300 miles? How long does it take for the freight train to go 220 miles? (These will be expressions in x.) Set those expressions equal to each other and solve for x.

OpenStudy (anonymous):

Thank you for your response. What I am having a hard time understanding is how to write the equations to get the answer. Can you give an example of one so I can finish off the rest? Thank you.

OpenStudy (anonymous):

Try doing the first part of what I wrote. I'll tell you if you get it right or wrong.

OpenStudy (anonymous):

300=16*x???

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

I think I got it confused with how to write an equation for a problem earlier on in the course. x_x (Anxiety)

OpenStudy (anonymous):

Call the speed of the freight train x. What's the speed of the passenger train? Hint: The answer is in the first sentence of the problem.

OpenStudy (anonymous):

300.

OpenStudy (anonymous):

300 doesn't appear in the first sentence of the problem. Explain in your own words what the first sentence says.

OpenStudy (anonymous):

16mph. (Oh my, please don't give up on me lol)

OpenStudy (anonymous):

I won't give up, as long as you try. Again, explain in your own words what the first sentence says.

OpenStudy (anonymous):

The passenger train is faster than the freight train...

OpenStudy (anonymous):

Good. How much faster?

OpenStudy (anonymous):

16mph

OpenStudy (anonymous):

So, as an example, if the freight train was going 50mph, how fast would the passenger train be going?

OpenStudy (anonymous):

66mph?

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

I added 16mph to the passenger train because it's faster. (I don't think that's how this is supposed to be solved though.)

OpenStudy (anonymous):

If the freight train was going x mph, how fast would the passenger train be going?

OpenStudy (anonymous):

300?

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

Oops. 16mph?

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

I apologize, I am having a hard time understanding this. I am no good at word problems but I will keep trying. D=RT As you stated above. 300=(r+16)T for the passenger train???

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

Distance=Rate*Time. So, I think this is the right equation for the passenger train. Is it right?

OpenStudy (anonymous):

First tell me how you got it.

OpenStudy (anonymous):

Distance is 300. Rate is 16mph * Time but we don't know what time is yet.

OpenStudy (anonymous):

Where did the r come from?

OpenStudy (anonymous):

I placed the r there for Rate.

OpenStudy (anonymous):

Rate of what?

OpenStudy (anonymous):

How fast the train was going.

OpenStudy (anonymous):

Which train? There are 2.

OpenStudy (anonymous):

the passenger train

OpenStudy (anonymous):

Why the passenger train?

OpenStudy (anonymous):

I think I created that equation first because it was the first train the question mentions. (That probably wasn't a good idea.)

OpenStudy (anonymous):

I never said anything was wrong. I just want to know how you got these answers to my questions. I can't help you understand the problem unless I know how you understand it. Now, in your equation, 300(r+16)t, what does r represent?

OpenStudy (anonymous):

Okay. So r represents the rate in the passenger train.

OpenStudy (anonymous):

Then how did you get r+16 in the equation?

OpenStudy (anonymous):

Hm. I think I meant to write it as a multiplication to multiply rate * time.

OpenStudy (anonymous):

Okay, let's go back to my example. If the freight train is going 50mpg, how fast is the passenger train going? The answer is in the first sentence of the question.

OpenStudy (anonymous):

The passenger train is going 16mph

OpenStudy (anonymous):

Not if the freight train is going 50mph. Again, explain what the first sentence of the problem is telling you.

OpenStudy (anonymous):

The passenger train is 16mph faster than the freight train.

OpenStudy (anonymous):

Right. So if the freight train is going 50mpg, how fast is the passenger train going?

OpenStudy (anonymous):

I am still thinking 66mph and that's where I got the +16 because the passenger train is going 16mph faster than the freight train.

OpenStudy (anonymous):

Right. So write a mathematical equation explaining exactly how you got that answer.

OpenStudy (anonymous):

The equation is simple. It's just arithmetic.

OpenStudy (anonymous):

Your answer was right. It shows the correct thinking. Now you just need to turn it into a mathematical expression.

OpenStudy (anonymous):

The equation is simple. It's just arithmetic.

OpenStudy (anonymous):

Was that the same mathematical expression I had posted earlier? 300=16/r for the passenger train so for the freight train, I came up with 220=(r+16)t. So 300/r=220/r+16. I got a little extra help from the text book. I believe these are correct since each train can travel for the same length of time.

OpenStudy (anonymous):

That's way ahead of where we are right now. Write a mathematical express that shows how you got 66mph as the answer to my question.

OpenStudy (anonymous):

Okay. 66=(r+16)t I got to this answer with help from the previous equation.

OpenStudy (anonymous):

I said if the freight train is going 50mpg, how fast is the passenger train going? You said "I am still thinking 66mph and that's where I got the +16 because the passenger train is going 16mph faster than the freight train." I want you to write a mathematical expression explaining how you got 66 as the answer to this question.

OpenStudy (anonymous):

Okay. 50=(r+16)t

OpenStudy (anonymous):

How did you solve the problem I asked?

OpenStudy (anonymous):

I added the rate at 16mph to the 50mph freight train.

OpenStudy (anonymous):

And what was the answer?

OpenStudy (anonymous):

... And the answer was 66.

OpenStudy (anonymous):

Right. Now write the mathematical expression you used to solve that problem.

OpenStudy (anonymous):

50+16=66.

OpenStudy (anonymous):

YES! Now if the freight train is going r mpg, how fast is the passenger train going?

OpenStudy (anonymous):

(Do I write a mathematical expression?) The passenger train is going 66mph.

OpenStudy (anonymous):

The 50/66 example was just an example to get you to understand what the first sentence is saying. Now I want you to write it as a more general expression, using a variable, r. We're going to call the speed of the freight train r. That way we can write an expression for the speed of the passenger train. So, using the first sentence, and assuming the speed of the freight train is r, how would you express the speed of the passenger train?

OpenStudy (anonymous):

I'm starting to get this a little more, I think. 300=(r+16)t?

OpenStudy (anonymous):

You're jumping way ahead. Just look at the first sentence. If we call the speed of the freight train r, what's an expression for the speed of the passenger train?

OpenStudy (anonymous):

300=16+r

OpenStudy (anonymous):

How did you get that from the first sentence?

OpenStudy (anonymous):

300 being from the speed of the passenger train= 16 is distance and r is for the freight train.

OpenStudy (anonymous):

The first sentence of the problem doesn't mention 300 or distance.

OpenStudy (anonymous):

Please type the first sentence of the problem, word-for-word.

OpenStudy (anonymous):

The speed of a passenger train is 16mph faster than the speed of a freight train. So 16+r?

OpenStudy (anonymous):

Yes! Do you see how that sentence correlates with that mathematical expression?

OpenStudy (anonymous):

I see. :)

OpenStudy (anonymous):

So that's where r+16 comes from.

OpenStudy (anonymous):

Okay. So now we have the speeds for the two trains: r for the freight train, and r+16 for the passenger train. Let's just work with the passenger train first. Using the formula distance = rate times time (d=rt), write an equation for how far the passenger train goes in time t. You can use data from the second sentence.

OpenStudy (anonymous):

t=300/r?

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

I got that by using the formula d=rt. However, I do not think this is gonna work. So, 300/r+16? or 300=(r+16)t?

OpenStudy (anonymous):

300=(r+16)t is right. But why, exactly?

OpenStudy (anonymous):

300 being the speed of the train. r is the rate + 16 * time.

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

Scratch the previous answer. The passenger train is going 16mph faster than the freight train. So that is r+16.

OpenStudy (anonymous):

Correct. So what's the d=rt equation for the passenger train? 1. What distance does it travel? 2. What's its rate (speed)? 3. How long does it take?

OpenStudy (anonymous):

300=(r+16)t. Okay.

OpenStudy (anonymous):

Good! Now write an equation for the freight train.

OpenStudy (anonymous):

220=(r+16)t. Distance traveled, rate (speed)* time

OpenStudy (anonymous):

Not exactly. What the speed of the freight train?

OpenStudy (anonymous):

D'oh! I think that in this case we do, 220/r because we do not have rate for the freight train. So the rate for now is r???

OpenStudy (anonymous):

Well, we just assigned the variable r to be the speed of the freight train. The we described the speed of the passenger train in relation to that variable, r. Exactly like the first sentence described. So, write out the equation for the freight train, just like you did for the passenger train. 1. What distance does it travel? 2. What's its rate (speed)? 3. How long does it take?

OpenStudy (anonymous):

220=(r+r)t??

OpenStudy (anonymous):

Where did you get r+r?

OpenStudy (anonymous):

I got confused. So going back three of your questions. The speed of the freight train is r. 220=(r)t?

OpenStudy (anonymous):

Yes. Now, before we go any further, look back at the first sentence of the problem and remind yourself of how we got r and r+16 for the speeds of the trains. Then look at the d=rt equations you wrote for the two trains, and remind yourself of how we got them from the information in the first two sentences.

OpenStudy (anonymous):

This is the gist of story problems. There are several quantities, and they tell you a story about them. Some quantities won't be known, so you'll have to pick variables to assign to them. Some quantities WILL be known (they'll just be numbers). And then they'll give you some ways to relate some of the quantities to some other quantities. You have to write those relationships as mathematical equations, so that you can solve them.

OpenStudy (anonymous):

Will do. Does it really matter when you choose any letter as a variable? Or can it be just any to represent that missing number? I took notes of this. I think I'm prepared to answer more on my own for this week's assignment thanks to you.

OpenStudy (anonymous):

In this problem, we're dealing with two speeds, two distances, and two times. We can use those to write two d=rt equations. They tell us one speed in terms of the other. They give us explicit values for the distances. And they tell us the times are equal. Once we write the two equations, we have to solve them together. That's the next step.

OpenStudy (anonymous):

Okay.